In the table below, exchange rate is defined as US dollars per Euro, E$/€. Given the information below, using UIP (in approximate form), fill in the blanks marked with letters. Round your answers to 3 decimals.

Interest Rate on Dollar Deposit (annual)

Interest Rate on Euro Deposit (annual)

Spot Exchange Rate, E$/€ (today)

Expected Future Exchange Rate, Ee$/€ (in one year)

Expected Rate of Change in Exchange Rate (Ee$/€-E$/€)/E$/€

Expected Dollar Return on Euro Deposit (annual)

Investors Prefer

0.025 (2.5%)

0.005 (0.5%)

1.10

1.326

0.205 (20.5%)

0.21

A

0.025

0.005 (0.5%)

B

1.326

0.105

C

D

0.025

0.005

E

1.326

F

G

Indifferent

0.025

0.005

1.4

1.326

H

I

J
Answer #1

1. As the dollar return on Euro Deposit is 0.21 or 21%, which is much higher than the interest rate in either Dollar or Euro, investors will prefer borrowing in Dollars and Investing in Euro (A)

2. (1.326-B)/B = 0.105

=> B= 1.326/1.105 = 1.2

Take 1 USD , get 1/1.2 = 0.8333 Euro, invest at 0.5%, get 0.8375 Euro after 1 year, convert to 1.110525 USD (at the expected exchange rate in one year). So, C = 1.110525-1 = 0.110525

As the dollar return on Euro Deposit is 0.1105 or 11.05%, which is much higher than the interest rate in either Dollar or Euro, investors will prefer borrowing in Dollars and Investing in Euro (D)

3. As the investors are indifferent,Interest rate parity holds

Spot exchange rate= expected exchange rate in one year * (1+ euro interest rate)/(1+ USD interest rate)

E =1.326 *1.005/1.025 = 1.300127

Rate of change in Exchange rate F  = (1.326 -1.300127)/1.300127 = 0.019900

Take 1 USD , get 1/1.300127 = 0.769156 Euro, invest at 0.5%, get 0.773002 Euro after 1 year, convert to 1.025 USD (at the expected exchange rate in one year). So, G = 1.025-1 = 0.025

4.Rate of change in Exchange rate H = (1.326-1.4)/1.4 = -0.05286

Take 1 USD , get 1/1.4 = 0.7142857 Euro, invest at 0.5%, get 0.717857 Euro after 1 year, convert to 0.951878 USD (at the expected exchange rate in one year). So, I = 0.951878-1 = - 0.04812

As the dollar return on Euro Deposit is -0.048 or -4.8%, which is much lower than the interest rate in either Dollar or Euro, investors will prefer borrowing in Euro and Investing in Dollars (J)